# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ Abstract
Most existing necessary conditions for optimal control based on adjoining methods require both state information and costate information, yet the lack of costates for a given feasible trajectory in practice impedes the determination of optimality. This paper establishes a novel theoretical framework for time-optimal control of controllable linear systems, proposing the augmented switching law that represents the input control and the feasibility in a compact form. Given a feasible trajectory, the disturbed trajectory under the constraints of augmented switching law is guaranteed to be feasible, resulting in a novel state-centric necessary condition without dependence on costate information. A first order necessary condition is proposed that the Jacobian matrix of the augmented switching law is not full row rank, which also results in an approach to optimizing a given feasible trajectory further. The proposed necessary condition is applied to the chain-of-integrators systems with full box constraints, contributing to some conclusions challenging to reason by traditional costate-based necessary conditions.
@@ I Introduction
Time-optimal control for controllable linear systems achieves universal applications in aerospace [1], manufacturing [2, 3], robotic control [4, 5], and autonomous driving [6]. Numerous works have been conducted on the behaviors of the optimal controls, resulting in some well-known necessary or sufficient conditions [7]. However, while referring to existing conclusions of optimal control for problems in general forms, it remains difficult to fully solve a high order problem in practice, especially when state inequality constraints are introduced into the system [8]. Specifically, when given a planned feasible trajectory, it is challenging to determine whether the trajectory is optimal based on state information since most existing necessary conditions require costate information.
@@
In the domain of optimal control theory, numerous necessary conditions are developed based on the variational method [9] and Pontryagin’s maximum principle (PMP) [10]. The former one provides a necessary condition on the extreme of a functional in an open feasible set, while PMP has the potential to deal with inequality constraints on states or controls [11]. Among them, the state-control inequality constraints pose a challenge in solution since they involve the connection of different arcs. Chang [12], Dreyfus [13], and other scholars developed the direct adjoining method for the connection of unconstrained arcs and constrained arcs, i.e., the junction conditions of costates. Jacobson et al. [14] pointed out that constrained arcs are not allowed when the state constraint is of odd order $p>1$, and instead, the unconstrained arc is tangent to the constrained boundary at a single point. Makowski and Neustadt [15] generalized the PMP-based necessary condition to the mixed constraints of state and control. Bryson et al. [16] developed the necessary condition based on the indirect adjoining method, further supplemented by Kreindler [17]. Hartl et al. [7] systematically reviewed existing PMP-based necessary conditions and summarized their general forms. The necessary conditions are usually applied to rule out some forms of non-optimal controls in practice. In some simple cases, the optimal control can be fully solved based on some necessary conditions [18, 19], while in other cases, the necessary conditions are applied to guide the design of suboptimal trajectories [8]. However, the above necessary conditions require information on the states and the costates simultaneously, while the costates are not directly observable in practice. Given a planned feasible trajectory, it is challenging to determine whether the trajectory is optimal based on the pure information on states, let alone the further optimization based on the given trajectory.
@@
In this paper, a necessary condition is called to be state-centric if it only requires information on states and controls, without dependence on costate information. There have existed some state-centric necessary so far, most of which are based on Bellman’s principle of optimality [20]. Bellman’s principle of optimality states that a sub-arc of an optimal trajectory is also optimal for the induced sub-problem, where the costates are not introduced. The Hamilton-Jacobi-Bellman (HJB) equation [21] represents a state-centric necessary in the form of the partial differential equation, which is widely used in optimal control [22] including reinforcement learning [23]. Evans and James [24] investigated the HJB equation for time-optimal control. Wolenski and Zhuang [25] showed that the minimal time function is the unique proximal solution to the HJB equation. However, most numerical methods based on the HJB equation require high computational cost [26], where the continuous time domain is discretized into a grid. A trade-off exists between the computational cost and the computational accuracy in practice. It remains challenging to develop a state-centric necessary condition for time-optimal control in a compact form with limited computational cost.
This paper sets out to establish a theoretical framework for time-optimal control of controllable linear systems and develop a novel state-centric necessary condition. The optimal control problem is formulated and investigated based on Hamiltonian in Section II. Section III analyzes the structure of arcs and the switching law of the optimal trajectory, where the connections between adjacent arcs are fully discussed. The keypoints for feasibility are introduced in Section IV, where the augmented switching law is proposed to represent the control and the feasibility in a compact form. Based on the uniqueness of the optimal control and the feasibility of the disturbed trajectory, Section V develops a novel state-centric necessary condition for time-optimal control. The proposed necessary condition is further applied to the chain-of-integrators systems with full box constraints in Section VI as examples. The contributions of this paper are as follows:
@@
    This paper establishes an innovative theoretical framework for time-optimal control of controllable linear systems, proposing the augmented switching law which fully represents the input control and the feasibility of the Bang-Singular-Bang trajectory in a compact form. The established framework focuses on the control of each unconstrained arc and constrained arc, the constraints induced by the connections of adjoining arcs, the feasibility of the disturbed trajectory near the constrained boundary, and the optimality of a given feasible trajectory. In the proposed theory, the control can be fully represented by the augmented switching law and the motion time of each arc, where the constraints directly induced by the augmented switching law can guarantee the feasibility of the trajectory near the constrained boundary in a sufficient and necessary sense. The developed framework provides a novel variational approach as well as a local optimization method with feasibility assurance, i.e., disturbing the motion time of each arc under fixed augmented switching law, resulting in disturbed trajectories with feasibility guarantees.
@@
    This paper proposes a novel state-centric necessary condition for optimal control of controllable linear systems, which only requires the state and control information without dependence on costate information. Given a Bang-Singular-Bang feasible trajectory which is claimed to be optimal, all disturbed feasible trajectories should result in a terminal time which is strictly longer than that of the original trajectory. In this way, a first order necessary condition is proposed that the Jacobian matrix of the augmented switching law should not be full row rank. The higher order necessary conditions are also expected to be defined based on the strict local optimality of the trajectory. Based on the proposed necessary condition, a novel optimization approach with a given Bang-Singular-Bang feasible trajectory as initial value is proposed, where the optimized trajectory always has the same augmented switching law as the initial trajectory in a Bang-Singular-Bang form and the feasibility is always guaranteed during iteration.
@@
    The proposed state-centric necessary condition is applied to high-order chain-of-integrators systems with full box constraints as examples, contributing to some conclusions that are challenging to prove by traditional costate-based necessary conditions, which demonstrates the effectiveness of the proposed necessary condition. Note that time-optimal control for chain-of-integrators systems with full box constraints represents a challenging and open problem in optimal control theory and kinematics, yet to be resolved. The number of arcs is proved to be less than the order of the system plus the sum of the order of arcs’ constraints under some conditions, while the traditional methods have to analyze the costates case-by-case. The proposed state-centric necessary condition successfully proves the recursive equations of motion time if the chattering phenomenon occurs and is induced by the 2nd order state constraints, which is challenging based on costate analysis due to the high degree-of-freedom of costates in high order problems. As a result, it is hopeful to determine the existence of chattering induced by the 2nd order state constraints in chain-of-integrators systems in future works.
II Problem Formulation and Assumptions
This section formulates the time-optimal control problem for controllable linear systems and introduces some necessary assumptions. The optimal control problem is defined as follows:
@@ II Problem Formulation and Assumptions
$\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right].$ (1e)
[…]
@@
Generally, an equality constraint ${\boldsymbol{F}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ can be introduced into problem (1a). However, it is not considered in this paper since the equality constraint can be equivalently eliminated by a linear transformation.
@@
Among them, $\lambda_{0}\geq 0$ is a constant. ${\boldsymbol{\lambda}}=\left(\lambda_{k}\right)_{k=1}^{n}$ is the costate vector. $\lambda_{0}$ and ${\boldsymbol{\lambda}}$ satisfy $\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right)\right)\not=0$. The Euler-Lagrange equations [27] holds that
@@
$\eta_{p}\geq 0,\,\eta_{k}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)=0,\,\forall 1\leq p\leq P.$ (5)
@@
$u\left(t\right)\in\left[-{u_{\mathrm{m}}},{u_{\mathrm{m}}}\right]$ is undetermined during ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$, which is called a singular arc. The notation $\equiv$” means that the equality holds for a continuous period. (7) is the well-known Bang-Singular-Bang law [8].
@@ III Arc Representation of the Optimal Trajectory
An arc of a trajectory is defined as a continuous segment of the optimal trajectory conditioned by the same input control. Two kinds of arcs exist, i.e., unconstrained arcs and constrained arcs, which are discussed in Section III-A and Section III-B, respectively. Through discussion on the two arcs, the system behavior representing a single arc is defined in Section III-C. Finally, if the chattering phenomenon does not occur, an optimal trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-D.
@@ III-A Unconstrained Arcs in Problem (1a)
In an unconstrained arc, ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ holds strictly almost everywhere, and the control does not switch. The following proposition states that $u\equiv\pm{u_{\mathrm{m}}}$ in an unconstrained arc.
@@
In an unconstrained arc, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not=0$ almost everywhere, and $u\equiv-{u_{\mathrm{m}}}{\mathrm{sgn}}\left({\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\right)$.
@@
Assume that a singular condition holds in the unconstrained arc, i.e., ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period. By (5), ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$ in an unconstrained arc. Hence, (4) implies that $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$. Considering the $k$-th order derivative of ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$, it has
$\forall k\in{\mathbb{N}},\,{\boldsymbol{\lambda}}^{\top}{\boldsymbol{A}}^{k}{\boldsymbol{b}}\equiv 0.$ (9)
By (2), ${\boldsymbol{\lambda}}\equiv 0$. Then, (8) implies that $\lambda_{0}=0$, which contradicts $\left(\lambda_{0},{\boldsymbol{\lambda}}\right)\not=0$. Therefore, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not\equiv 0$.∎

### v2

@@ Abstract
Most existing necessary conditions for optimal control based on adjoining methods require both state and costate information, yet the unobservability of costates for a given feasible trajectory impedes the determination of optimality in practice. This paper establishes a novel theoretical framework for time-optimal control of controllable linear systems with a single input, proposing the augmented switching law (ASL) that represents the input control and the feasibility in a compact form. Given a feasible trajectory, the perturbed trajectory under the constraints of ASL is guaranteed to be feasible, resulting in a novel state-centric necessary condition without dependence on costate information. A first-order necessary condition is proposed that the Jacobian matrix of the ASL is not full row rank, which also results in a potential approach to optimizing a given feasible trajectory with the preservation of arc structures. The proposed necessary condition is applied to high-order chain-of-integrator systems with full box constraints, contributing to some theoretical results challenging to reason by costate-based conditions.
@@ I Introduction
Time-optimal control for controllable linear systems achieves universal applications in aerospace [1], manufacturing [2, 3], robotic control [4, 5], and autonomous driving [6]. Numerous works have been conducted on the behaviors of the optimal controls, resulting in some well-known necessary or sufficient conditions [7]. However, it remains difficult to fully solve a high-order problem in practice, especially when state inequality constraints are introduced [8, 9]. Specifically, when given a feasible trajectory, it is challenging to determine the trajectory’s optimality based on state information since most existing necessary conditions require costate information, especially when the trajectory is not planned by costate-based methods.
@@
In the domain of optimal control theory, numerous necessary conditions are developed based on the variational method [10] and Pontryagin’s maximum principle (PMP) [11]. The former one provides a necessary condition on the extreme of a functional in an open feasible set, while PMP has the potential to deal with inequality constraints on states or controls [12]. Among them, the state-control inequality constraints pose a challenge in solution since they involve the connection of different arcs. Chang [13], Dreyfus [14], and other scholars developed the direct adjoining method for the connection of unconstrained arcs and constrained arcs, i.e., the junction conditions of costates. Jacobson et al. [9] pointed out that constrained arcs are not allowed when the state constraint is of odd order $p>1$, and instead, the unconstrained arc is tangent to the constraint’s boundary at a single point. Makowski and Neustadt [12] generalized the PMP-based necessary condition to the mixed constraints of state and control. Bryson et al. [15] developed the necessary condition based on the indirect adjoining method, further supplemented by Kreindler [16]. The necessary conditions are usually applied to rule out some forms of non-optimal controls in practice. However, the above necessary conditions require information on both states and costates, while costates are usually unobservable in practice, resulting in a challenge in determining the optimality of a given feasible trajectory by costate-based conditions.
Based on PMP, one can get rid of dependence on costates by deriving the optimal trajectory’s structure. Switching surfaces can be constructed in the state space for time-optimal control of linear systems. Walther et al. [17] computated switching surfaces for some systems through Gröebner basis. Patil et al. [18] represented switching surfaces using state equations for single-input diagonal linear systems without state constraints. However, for systems with state constraints, it remains challenging to completely construct switching surfaces due to the lack of optimal switching laws. For example, He et al. [19] provided explicit equations for switching surfaces of 3rd-order chain-of-integrator systems with full state constraints, limited to zero terminal states. Wang et al. [20] extended [19] to non-zero terminal conditions, failing to achieve time-optimality for higher-order systems. Numerical methods [21] face the curse of dimensionality. Hence, there is a lack of efficient and general optimality conditions without dependence on costates for high-order systems with state constraints.
@@
Another type of necessary condition leverages Bellman’s principle of optimality [22], asserting that a sub-arc of an optimal trajectory is itself optimal for the induced sub-problem, without explicitly involving costates. The Hamilton-Jacobi-Bellman (HJB) equation [23] represents a typical necessary condition in the form of partial differential equations, which is widely used in optimal control [24, 25]. Evans and James [26] investigated the HJB equation for time-optimal control. Wolenski and Zhuang [27] showed that the minimal time function is the unique proximal solution to the HJB equation. However, most numerical HJB-based methods require high computational cost [28] due to the discretized grids. From a theoretical perspective, HJB-based methods also fail to preserve arc structures, which hinders the analytical reasoning.
Compared with the PMP-based and HJB-based necessary conditions, this paper sets out to establish a theoretical framework for time-optimal control of controllable single-input linear systems and develops a novel state-centric necessary condition that requires state information without dependence on costate information. Sections III and IV derive trajectories’ structures on arcs and keypoints for feasibility, respectively. Section V develops the state-centric necessary condition which is applied from theoretical and numerical perspectives in Sections VI and VII as examples, respectively. The contributions of this paper are as follows:
@@
    This paper establishes a novel theoretical framework for time-optimal control of controllable single-input linear systems, proposing the augmented switching law (ASL) which represents structures and feasibility of bang-bang and singular (BBS) trajectories in a compact form. Compared to existing representations of switching surfaces, the ASL can generally deal with high-order systems with state constraints. The developed framework provides a novel variational approach with feasibility assurance, i.e., perturbing the motion time of each arc with a fixed ASL and resulting in feasible perturbed BBS trajectories.
@@
    Based on the developed framework, this paper proposes a novel state-centric necessary condition for time-optimality, asserting that the Jacobian matrix induced by the ASL is not full row rank. The condition is derived from the fact that if a BBS trajectory is optimal, then all perturbed feasible trajectories have a strictly longer terminal time than the original one. Compared to existing costate-based conditions, the proposed state-centric necessary condition only requires state information without dependence on costates. Note that costates are usually unobservable in practice, especially when the trajectory is not planned by costate-based optimization.
@@
    This paper proposes a potential optimization idea based on the developed necessary condition. Given a feasible BBS trajectory as the initial value, the motion time is optimized with a fixed ASL until the Jacobian matrix is not full row rank. During iteration, the feasibility is guaranteed based on the developed ASL framework. Compared to discretized methods, the proposed approach can preserve the BBS arc structure and has the potential to significantly reduce control oscillations and open-loop integration errors of terminal states.
  4. 4.
    The proposed state-centric necessary condition is applied to high-order chain-of-integrator systems with full state constraints which remains an open problem in the optimal control domain. An upper bound on the number of arcs is determined in a general sense, while case-by-case analysis is necessary in traditional methods. A recursive equation of junction time is reasoned in chattering induced by 2nd-order constraints, which is challenging to prove by costate analysis. As a result, one could determine the existence of chattering induced by 2nd-order constraints in chain-of-integrator systems in future works. The above results demonstrate the theoretical effectiveness of the proposed necessary condition.
II Problem Formulation
Some notations should be introduced. For an integer $n\in{\mathbb{N}}$, the set $\left[n\right]$ refers to $\left\{1,2,\dots,n\right\}$. $\left(x_{i}\right)_{i\in\mathcal{I}}\in{\mathbb{R}}^{\left|\mathcal{I}\right|}$ is a column vector and $\left(x_{ij}\right)_{i\in\mathcal{I},j\in\mathcal{J}}\in{\mathbb{R}}^{\left|\mathcal{I}\right|\times\left|\mathcal{J}\right|}$ is a matrix, where $\mathcal{I}$ and $\mathcal{J}$ are index sets. Specifically, $\left(x_{i}\right)_{i=1}^{n}\in{\mathbb{R}}^{n}$ refers to $\left(x_{i}\right)_{i\in\left[n\right]}$.
This section formulates the time-optimal control problem for controllable linear systems and introduces some necessary assumptions. The optimal control problem is as follows:
@@ II Problem Formulation
$\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (1e)
where ${\boldsymbol{A}}\in{\mathbb{R}}^{n\times n}$, ${\boldsymbol{b}}\in{\mathbb{R}}^{n}$, ${\boldsymbol{C}}=\left({\boldsymbol{c}}_{p}\right)_{p=1}^{P}\in{\mathbb{R}}^{P\times n}$, and ${\boldsymbol{d}}=\left(d_{p}\right)_{p=1}^{P}\in{\mathbb{R}}^{P}$, where $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$. In problem (1a), the terminal time ${t_{\mathrm{f}}}$ is free. ${\boldsymbol{x}}=\left(x_{k}\right)_{k=1}^{n}\in{\mathbb{R}}^{n}$ is the state vector. The state constraint (1c) requires that $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$. The input control $u$ is subject to a box constraint (1e).
@@
Generally, an equality constraint ${\boldsymbol{F}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ can be incorporated into problem (1a). However, it is not considered in this paper since the equality constraint can be equivalently eliminated by a linear transformation.
@@
Among them, $\lambda_{0}\geq 0$ is a constant. ${\boldsymbol{\lambda}}=\left(\lambda_{k}\right)_{k=1}^{n}$ is the costate vector. $\forall t\in\left[0,{t_{\mathrm{f}}}\right]$, $\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right)\right)\in{\mathbb{R}}^{n+1}$ is not ${\boldsymbol{0}}$ [8]. The Euler-Lagrange equations [8] holds that
@@
$\eta_{p}\geq 0,\,\eta_{p}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)=0,\,\forall p\in\left[P\right].$ (5)
@@
$u\left(t\right)\in\left[-{u_{\mathrm{m}}},{u_{\mathrm{m}}}\right]$ is undetermined during ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$, which is called a singular arc. The notation $\equiv$” means that the equality holds for a continuous period. (7) is the well-known bang-bang and singular (BBS) control law [11, 29]. In other words, the optimal control is always ±$\pm{u_{\mathrm{m}}}$, except during singular arcs.
@@
Given a feasible trajectory $\left({\boldsymbol{x}}\left(t\right),u\left(t\right),\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right)\right)$, one can determine the optimality through PMP-based necessary conditions like (4)–(8). However, in practice, only $\left({\boldsymbol{x}}\left(t\right),u\left(t\right)\right)$ is observable, while $\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right)\right)$ is difficult to fully solve based on provided $\left({\boldsymbol{x}}\left(t\right),u\left(t\right)\right)$. Consequently, the focus of this paper is to determine the optimality of a feasible trajectory in problem (1a) using the provided $\left({\boldsymbol{x}}\left(t\right),u\left(t\right)\right)$.
@@ III Arc Representation of the Optimal Trajectory
An arc of a trajectory is defined as a continuous segment conditioned by a fixed dynamic equation, i.e., $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ where $\widehat{{\boldsymbol{A}}}$ and $\widehat{{\boldsymbol{b}}}$ are constant. Two kinds of arcs, i.e., unconstrained arcs and constrained arcs, are discussed in Section III-A and Section III-B, respectively. Through discussion on the two arcs, the system behavior representing a single arc is defined in Section III-C. Finally, an optimal non-chattering trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-D.
@@ III-A Unconstrained Arcs in Problem (1a)
In an unconstrained arc, ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ holds almost everywhere (a.e.), and the control does not switch. The following proposition states that $u\equiv\pm{u_{\mathrm{m}}}$ in an unconstrained arc.
@@
In an unconstrained arc, it holds a.e. that ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not=0$ and $u\equiv-{u_{\mathrm{m}}}{\mathrm{sgn}}\left({\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\right)$.
@@
Assume that a singular condition holds in the unconstrained arc, i.e., ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period. By (5), ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ implies that ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$ holds a.e. in an unconstrained arc. Hence, (4) implies that $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$. Therefore,
$\forall k\in{\mathbb{N}},\,{\boldsymbol{\lambda}}^{\top}{\boldsymbol{A}}^{k}{\boldsymbol{b}}=\frac{\mathrm{d}^{k}}{\mathrm{d}t^{k}}\left({\boldsymbol{\lambda}}^{\top}{\boldsymbol{b}}\right)\equiv 0,$ (9)
i.e., ${\boldsymbol{\lambda}}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\equiv{\boldsymbol{0}}$. However, by (2), ${\boldsymbol{\lambda}}\equiv 0$ holds a.e. Then, (8) implies that $\lambda_{0}=0$, which contradicts $\left(\lambda_{0},{\boldsymbol{\lambda}}\right)\not=0$. Therefore, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not\equiv 0$ holds a.e. ∎